# Random time-shift approximation enables hierarchical Bayesian inference of mechanistic within-host viral dynamics models on large datasets

### v1

@@ 3.2 NBA study
In the main text we report only the marginal posteriors for the shared parameters and posterior predictive VL trajectories for select individuals. The marginal posteriors for the shared parameters are shown in Fig. 9. We see that there is large movement in the distributions from the prior (orange) to the posteriors (blue) for all parameters besides $\sigma_{R_{0}}$. This is not surprising since the individual parameters $R_{0}$, $\delta$, and $\rho$ values trade-off with one another and with less data in the growth phase, this can cause the related hyper-parameters to be weakly identifiable when inferred jointly. The model appears to be able to estimate all parameters including the measurement noise $\kappa$. Our estimates for $\mu_{R_{0}}$—and by extension—the individual $R_{0}$’s are consistent with the results in Table 2. It should be noted that compared to Zitzmann et al. [5] our population estimate is larger $15.3$ compared with $8.2$. One possible reason for this is that our model accounts for the probability of extinction during the early phase which is non-negligible when assuming the infection is initiated by a single cell in the eclipse phase. This potentially biases the results in [5] downwards. Our $R_{0}$ estimates are consistent with the value of 14.2 reported in Germano et al. [18] from their stochastic model (which also accounts for immune responses). Our population estimate for $\mu_{\rho}$ is $6.57$ which is larger than the estimate of $3.07$ in [5]. We infer the population mean rate of removal for infectious cells to be $1.27$ as in Zitzmann et al. [5]. There are several possible reasons for the difference in the estimates of $R_{0}$ and $\rho$. The first being that, as our model is stochastic, larger values of the parameters are required to ensure that the model overcomes extinction than what would be estimated by a deterministic model. We are also fitting to a much larger number of individuals (163 vs 25) including those that are unvaccinated and have limited data during the growth and decline phases. Additionally, we account for the LOD in our model and make no assumptions as to the time from infection to peak instead allowing our model to infer the infection time.
{adjustwidth}
-2in0in Refer to caption
@@
Fig. 10 shows the posterior predictive VL fits (blue) for a subset of individuals in the NBA dataset. The first row of Fig. 10 corresponds to data for four individuals also analysed in Zitzmann et al. [5] alongside their estimated VL trajectories (orange crosses). We see strong agreement between our model and the Zitzmann et al. [5] model, with our median predictions closely tracking their estimated trajectories.
The second row of plots corresponds to individuals (visually) selected for having limited data in either the pre- or post-peak phase. The posterior predictive trajectories generally track the observed data well, though some interesting features emerge for individuals 2987 and 10981. For these two, the light-blue prediction intervals indicate the possibility of earlier infection times. Given the limited data during the increasing or declining phases of their infections, this behaviour is plausible.
{adjustwidth}
-2.0in0in Refer to caption
Fig 10: Posterior predictive fits for a subset of individuals from the NBA dataset. The first row of plots corresponds to a subset of individuals also presented in Zitzmann et al. [5] for comparison purposes. The second row corresponds to individuals not studied in Zitzmann et al. [5], but that have sparse data (which were selected by visual inspection). The observations for each individual are shown in black. The solid lines represent the median of the posterior predictive fits and the shaded regions represent the 80% and 95% prediction intervals. The orange crosses in the first row of plots indicate the deterministic VL trajectories estimated in [5]. All subplots show the same time period $[-10,21]$.
@@ 4 Discussion
We have presented a novel method for fitting mechanistic within-host viral dynamics models that explicitly account for process noise, providing a more realistic representation of viral load (VL) trajectories. Our approach leverages the concept of a random time-shift—first introduced in Barbour et al. [16] and later developed into a practical tool in Morris et al. [17]—to capture the effects of stochasticity in the WHVDs. A key contribution of this work is demonstrating how the time-shift distributions can be directly incorporated into the likelihood function, allowing stochasticity to be modelled without the need for computationally expensive stochastic simulations, for example using Gillespie’s algorithm [22] or tau-leaping [23]. This enables efficient inference, making it feasible to fit the model to large datasets using a hierarchical framework. Our approach can be viewed as a hybrid method, combining the interpretability and tractability of deterministic models while capturing key elements of stochastic modelling, offering a middle ground between fully stochastic simulation and purely deterministic fitting.
Hierarchical approaches have been widely used in phenomenological studies of WHVDs [6, 11], but to our knowledge, no stochastic WHVD models have incorporated hierarchical structures. The existing approaches fall into two main categories: simpler parametric models, such as piecewise linear functions [4, 11, 6], and mechanistic models that assume deterministic dynamics [12, 5, 38, 25]. Parametric models offer computational advantages, as their likelihoods are easier to evaluate and they integrate well with probabilistic programming frameworks like Stan [56]. However, they primarily describe viral dynamics in terms of slope parameters (for growth and decline) and a hinge point (for peak VL), limiting their ability to extract biologically meaningful quantities like $R_{0}$ and production rates $\rho$. Mechanistic models, in contrast, provide a more realistic representation of infection dynamics but are computationally expensive due to the need to repeatedly solve ODEs. This challenge is especially pronounced in Bayesian inference for large datasets. Moreover, parameter identifiability can become a major concern in this context, particularly when biological parameters influence the VL in overlapping ways (i.e. the relationship between $R_{0}$, the viral production rate, the infection time and initial conditions) [5, 25]. While some hybrid stochastic-deterministic models [57, 58, 18] have been developed, they remain computationally demanding and have not been widely applied to multi-individual VL kinetics studies. Our method addresses these limitations by introducing stochasticity directly into the model formulation, capturing process noise without the need for costly stochastic simulations. This ensures that parameters are not artificially inflated while maintaining computational efficiency.
@@
The primary limitation of our method is its complexity compared to simpler approaches in the literature that leverage standard packages like Stan [56]. Our framework integrates multiple techniques—neural networks, random time-shift computations, Laplace approximations, and numerical solvers—to fit the model. The neural network plays a critical role in enabling the time-shift approximation by learning a smooth mapping from model parameters to the parameters of the time-shift distribution, effectively circumventing the need for expensive optimisation routines during likelihood evaluation. This learned surrogate is then embedded directly within the likelihood, facilitating fast and efficient likelihood computations. This complexity means the method is not a black box and requires expertise to debug. If performance issues arise, such as MCMC convergence problems, identifying the source of the issue within the pipeline can be challenging.
@@
Increased method complexity often comes with computational challenges, akin to the trade-offs between importance sampling in particle filters and bootstrap filtering methods [61, 62]. To mitigate these issues, we have structured the method into several modular and easily testable steps. The two main approximations—using a neural network to learn time-shift distribution parameters and employing Laplace approximations—can both be replaced with exact computations (at the cost of efficiency) for validation. We recommend validating each component separately before applying the full framework. This includes ensuring that the neural network generalizes well beyond its training set, verifying the accuracy of time-shift distributions (e.g., by comparing against simulations from the stochastic model [17]), and using Laplace approximations only when justified (e.g., for normally distributed observation processes). Profile likelihoods can further confirm that the method accurately captures the shape of the exact likelihood, as demonstrated in Section 4 of S1 text.
@@
Another consideration is that the branching process results for each model must be computed from scratch. Since time-shift distributions are model-specific, they must be re-derived, using the framework from Morris et al. [17]. Our previous work [17] provides a package for computing the time-shift distributions given the branching process results. Additionally, the neural network must be redesigned and retrained for any new VL kinetics model as the underlying branching process structure is altered. While these challenges are not insurmountable, they are important factors when deciding whether to use this approach. However, our work here provides a general template for implementation making it substantially easier to extend the method to other models going forward.
@@
A key advantage of our approach is its ability to model heterogeneity in viral trajectories, which is particularly relevant when initial infections involve a small number of cells. In such cases, accounting for the probability of early extinction may be important, as the likelihood of extinction is higher and could influence parameter estimates if unaccounted for. Beyond within-host modelling, incorporating stochasticity also has implications for between-host transmission studies [63, 64, 65, 66]. Small variations in VL trajectories can influence an individual’s infectiousness, affecting transmission dynamics in settings like household studies [37, 67, 68]. A more realistic representation of within-host variability could lead to improved estimates of key epidemiological parameters, which is an area of ongoing research.
@@
This paper introduces a methodological advance for working with within-host viral dynamics (WHVDs), addressing a complex and increasingly important class of inference problems. Rather than proposing a new model, we focus on developing computational tools that enable efficient Bayesian inference for existing stochastic VL kinetics models. Our interdisciplinary approach—combining applied mathematics, machine learning, and computational statistics—makes it feasible to fit these models to large datasets using biologically motivated, mechanistic, models. By leveraging modern computational techniques, we present a flexible framework that can accommodate the intricacies of viral dynamics and scale efficiently. A key enabler of this work is the use of the Julia programming language [41], which allows for seamless integration of model specification, automatic differentiation, and inference within a single environment. This capability makes it possible to conduct complex analyses on local hardware, reducing the reliance on dedicated high-performance computing clusters. Overall, the methods introduced here offer a powerful toolset for advancing the study of WHVDs and supporting future research into more realistic and scalable modelling frameworks.
@@
The following is the supporting information for the main text. It has been slightly adjusted to allow inclusion in a single preprint containing all materials.
S1 text
1. The deterministic approximation
Let $\boldsymbol{x}(t)=K^{-1}\boldsymbol{X}(t)$ denote the density process of the CTMC $\{\boldsymbol{X}(t),t\geq t_{0}\}$. Then reading down column 2 of Table 1 in the main text, and provided $\boldsymbol{a}+\boldsymbol{l}\in\mathcal{S}$ where $\mathcal{S}$ is the state space of the CTMC, $\boldsymbol{a}$ is the current state and $\boldsymbol{l}$ represents the jumps of the process, then the (positive) transition rates can be expressed in terms of the following functions
$r^{(K)}(\boldsymbol{a},\boldsymbol{a}+\boldsymbol{l})=\begin{cases}K\alpha sv,\quad&\textrm{if }\boldsymbol{\ell}=(-1,1,0,0),\\ Kke,\quad&\textrm{if }\boldsymbol{\ell}=(0,-1,1,0),\\ K\delta i,\quad&\textrm{if }\boldsymbol{\ell}=(0,0,-1,0),\\ K\rho i,\quad&\textrm{if }\boldsymbol{\ell}=(0,0,0,1),\\ Kcv,\quad&\textrm{if }\boldsymbol{\ell}=(0,0,0,-1).\end{cases}$ (18)
Since the rates can be expressed in this form (referred to as “density dependent”), by Theorem 3.1 of [35], as $K\to\infty$ the process $\boldsymbol{x}(t)$ converges to a deterministic process $\boldsymbol{x}_{d}(t)$ uniformly in probability over finite time intervals provided $\boldsymbol{x}_{d}(0)=K^{-1}\boldsymbol{X}(0)$. Letting
$\boldsymbol{H}(\boldsymbol{x}_{d}(t))=\sum_{\boldsymbol{l}}\boldsymbol{l}r(\boldsymbol{a},\boldsymbol{a}+\boldsymbol{l}),$
where $r(\boldsymbol{a},\boldsymbol{a}+\boldsymbol{l})=K^{-1}r^{(A)}(\boldsymbol{a},\boldsymbol{a}+\boldsymbol{l})$, then the system of ODEs governing the deterministic approximation (Eq. 2 in the main text) is given by
[…]
2. Branching process results
Considering the early stages of infection within a host, the number of target cells, $S$, is approximately constant, i.e. $S\approx S_{0}$ and hence the rates in Table 1 of the main text are linear. The BPA then tracks the evolution of the state vector $\boldsymbol{X}_{b}(t)=(E,I,V)$ and the change in event and corresponding rates are given in Table 4.
$\Delta\boldsymbol{X}$ Rate
$(1,0,0)$ $\beta S_{0}V$
$(-1,1,0)$ $kE$
$(0,-1,0)$ $\delta I$
$(0,0,1)$ $\rho I$
$(0,0,-1)$ $cV$
Table 4: Change in state and rates for the BPA to the $TCL$ model.
To apply the method in [17], we must identify the constant terms from the progeny generating functions which are defined as
$f_{i}(\boldsymbol{s})=\sum_{\boldsymbol{k}}p_{i}(\boldsymbol{k})\prod_{j=1}^{m}s_{j}^{k_{j}},\quad\boldsymbol{s}\in[0,1]^{m}$
where $p_{i}(\boldsymbol{k})$ is the probability of an individual of type $i$ having $\boldsymbol{k}=(k_{1},\dots,k_{m})$ offspring of each type [17]. Let $\beta^{\star}=\beta S_{0}$, then for the TCL model
$f_{1}(\boldsymbol{s})$ $=\frac{k}{k}s_{2}=s_{2}$
$f_{2}(\boldsymbol{s})$ $=\frac{\delta+\rho s_{2}s_{3}}{\delta+\rho}$
$f_{3}(\boldsymbol{s})$ $=\frac{c+\beta^{\star}s_{1}s_{3}}{c+\beta^{\star}}.$
The non-zero constants (see [17] for details) needed for the time-shift computations are
$\alpha_{12}=k,\quad\beta_{223}=\rho,\quad\beta_{313}=\beta^{\star}.$
The probability of ultimate extinction $q^{\star}$ is determined by solving for the minimal non-negative solution of the following system [69]
$f_{i}(\boldsymbol{q})=q_{i}\textrm{ for }i=1,2,3$
where the $f_{i}(\boldsymbol{q})$ are the progeny generating functions. Making the substitution and rearranging the system yields the following system of equations
$\rho(c+\beta)q_{3}^{2}-(c\delta+\rho\beta)q_{3}+c(\delta+\rho)$ $=0$ (20)
$q_{1}$ $=\frac{\delta}{\delta+\rho-\rho q_{3}}$ (21)
$q_{2}$ $=q_{1}.$ (22)
Since Eq. (20) is quadratic and only in terms of $q_{3}$, this can be solved analytically and the remaining probabilities determined by substitution.
3. Neural network architecture and details
When generating the training data, we follow the process outlined in Section 2.8.2 of the main text. One key detail from an implementation perspective is that we also standardise the inputs (model parameters) and the outputs (time-shift parameters) using standard scaling (i.e. subtract the sample mean for each parameter and divide by the sample variance). In the context of training NN’s, this standardisation has the effect of drastically improving the performance and convergence behaviour of optimisers like Adam. We utilise a simple NN with the following architecture:
  1. 1.
    Input layer: 3 neurons (the free parameters of the VL model) with ReLU activation
  2. 2.
    Hidden layer: 64 neurons with ReLU activation
  3. 3.
    Output layer: 3 neurons (the parameters of the time-shift distribution) with soft-plus activation
The input and hidden layer both use the rectified linear unit (ReLU) activation functions. The first of which acts to standardise the input data and then in the output layer we use the soft-plus activation function to ensure non-negativity in the GG parameters. We consider mean-squared-error (MSE) as the loss function to be optimised as this performed well in testing and all the output parameters for the surrogate distribution are on the same scale. Let
$\mathcal{A}=\left\{(\boldsymbol{\theta}_{i},\boldsymbol{\zeta}_{i})\right\}_{i=1}^{n_{\textrm{train}}}\quad\textrm{and}\quad\mathcal{B}=\left\{(\boldsymbol{\theta}_{i},\boldsymbol{\zeta}_{i})\right\}_{i=1}^{n_{\textrm{test}}}$
denote the training and validation sets respectively. Let $\mathbb{M}_{\Psi}(\boldsymbol{\theta})$ denote the current NN model—where we have collected the NN model parameters into $\Psi$—evaluated at mechanistic model parameters $\boldsymbol{\theta}$ which returns the surrogate distribution parameters. Define $\mathbb{M}(\boldsymbol{\theta})_{j}$ as the $j$th element of this (i.e. the estimated surrogate distribution parameter using the NN), then the loss is given by
$L(\Psi)=\sum_{i=1}^{n_{\textrm{train}}}\sum_{j=1}^{3}\left(\mathbb{M}_{\Psi}(\boldsymbol{\theta_{i}})_{j}-\zeta_{j}\right)^{2}.$ (23)
We use the ADAM optimiser with defaults as provided by the Julia package Flux.jl. Additionally to improve the generalisability of the model, we perform early stopping. Early stopping uses the validation set and we stop training the model if there is no improvement in the loss (calculated in the same way as the loss in Eq. (23)) on the validation set after 30 epochs. The model is completely implemented and trained in Julia on a CPU (to simplify dissemination of codes) using the Flux.jl package.
4. Comparison of our Laplace approximation versus the exact calculation
Figure 11 shows the profile likelihoods for a given parameter set (for a random individual). These fix all but the parameter of interest and allow that to vary over some interval. The orange dots show our approximation and we can see strong agreement between the approximation and the “exact” likelihood. The maximum (absolute) error computed over the points shown in the plot is $<1e-7$.
{adjustwidth}
-1in0in Refer to caption
Fig 11: Profile likelihoods for the exact method and the Laplace approximation. Profile likelihoods for the two methods: exact given by the solid blue line, Laplace approximation in orange points.

### v2

@@ 3.2 NBA study
The marginal posteriors for the shared parameters are shown in Fig. 6. We see that there is large movement in the distributions from the prior (orange) to the posteriors (blue) for all parameters besides $\sigma_{R_{0}}$. This is not surprising since the individual parameters $R_{0}$, $\delta$, and $\rho$ values trade-off with one another and with less data in the growth phase, this can cause the related hyper-parameters to be weakly identifiable when inferred jointly. The model appears to be able to estimate all parameters including the measurement noise $\kappa$. Our estimates for $\mu_{R_{0}}$—and by extension—the individual $R_{0}$’s are consistent with the results in Table 2. It should be noted that compared to Zitzmann et al. [5, SI] our population estimate is larger, $13.4$ compared with $8.2$. Our $R_{0}$ estimates are consistent with the value of 14.2 reported in Germano et al. [18] from their stochastic model (which also accounts for immune responses). Our population estimate for $\mu_{\rho}$ is $3.27$ which is slightly larger than the estimate of $3.07$, but still consistent with [5, SI]. We infer the population mean rate of removal for infectious cells to be $1.18$, slightly lower than the $1.3$ [5, SI]. There are several possible reasons for the difference in the estimates of $R_{0}$ (i.e. $\mu_{R_{0}}$). The first being that, as our model is stochastic, larger values of the parameters are required to ensure that the model overcomes extinction than what would be estimated by a deterministic model. We are also fitting to a much larger number of individuals (163 vs 25) including those that are unvaccinated and have limited data during the growth and decline phases. Additionally, we account for the LOD in our model and make no assumptions as to the time from infection to peak instead allowing our model to infer the infection time.
@@
Fig. 7 shows the posterior predictive VL fits (blue) for a subset of individuals in the NBA dataset. The first row of Fig. 7 corresponds to data for four individuals also analysed in the SI of Zitzmann et al. [5] alongside their estimated VL trajectories (red dashed line). We see strong agreement between our model and the model in the SI of Zitzmann et al. [5, SI] model, with our median predictions closely tracking their estimated trajectories. The second row of plots corresponds to individuals (visually) selected for having limited data in either the pre- or post-peak phase. The posterior predictive trajectories track the observed data well.
Fig 7: Posterior predictive fits for a subset of individuals from the NBA dataset. The first row of plots corresponds to a subset of individuals also presented in the SI of Zitzmann et al. [5] for comparison purposes. The second row corresponds to individuals not studied in that work, but that have sparse data that were selected by visual inspection. The observations for each individual are shown in black. The solid lines represent the median of the posterior predictive fits and the shaded regions represent the 80% and 95% prediction intervals. The red dashed lines in the first row of plots indicate the deterministic VL trajectories estimated in [5]. All subplots show the same time period [10,21][-10,21].
@@ 4 Discussion
We have presented a novel method for fitting mechanistic within-host viral dynamics models that explicitly account for process noise, providing a more realistic representation of viral load (VL) trajectories. Our approach leverages the concept of a random time-shift—first introduced in Barbour et al. [16] and later developed into a practical tool in Morris et al. [17]—to capture the effects of stochasticity in the WHVDs. Our key contribution in this work is demonstrating how the time-shift distributions can be directly incorporated into the likelihood function, allowing stochasticity to be modelled without the need for computationally expensive stochastic simulations, for example using Gillespie’s algorithm [22] or tau-leaping [23]. Furthermore, this enables efficient inference, making it feasible to fit the model to large datasets using a hierarchical framework. Our approach can be viewed as a hybrid method, combining the interpretability and tractability of deterministic models while capturing key elements of stochastic modelling, offering a middle ground between fully stochastic simulation and purely deterministic fitting.
Hierarchical approaches have been widely used in phenomenological studies of WHVDs [6, 11], but to our knowledge, no stochastic WHVD models have incorporated hierarchical structures. The existing approaches fall into two main categories: simpler parametric models, such as piecewise linear functions [4, 11, 6], and mechanistic models that assume deterministic dynamics [12, 5, 40, 25]. Parametric models offer computational advantages, as their likelihoods are easier to evaluate and they integrate well with probabilistic programming frameworks like Stan [55]. However, they primarily describe viral dynamics in terms of slope parameters (for growth and decline) and a hinge point (for peak VL), limiting their ability to extract biologically meaningful quantities like $R_{0}$ and production rates $\rho$. Mechanistic models, in contrast, provide a more realistic representation of infection dynamics but are computationally expensive due to the need to repeatedly solve ODEs. This challenge is especially pronounced in Bayesian inference for large datasets. Moreover, parameter identifiability can become a major concern in this context, particularly when biological parameters influence the VL in overlapping ways (i.e. the relationship between $R_{0}$, the viral production rate, the infection time and initial conditions) [5, 25]. While some hybrid stochastic-deterministic models [56, 57, 18] have been developed, they remain computationally demanding and have not been widely applied to multi-individual VL kinetics studies. Our method is the first to use a mechanistic model within a hierarchical framework that can also scale to incorporate hundreds of individuals.
@@
The primary limitation of our method is its complexity compared to simpler approaches in the literature that leverage standard packages like Stan [55]. Our framework integrates multiple techniques—neural networks, random time-shift computations, Laplace approximations, and numerical solvers—to fit the model. The neural network plays a critical role in enabling the time-shift approximation by learning a smooth mapping from model parameters to the parameters of the time-shift distribution, effectively circumventing the need for expensive optimisation routines during likelihood evaluation. This learned surrogate is then embedded directly within the likelihood, facilitating fast likelihood computations. This complexity means the method is not a black box and requires expertise to debug. If performance issues arise, such as MCMC convergence problems, identifying the source of the issue within the pipeline can be challenging.
@@
Increased method complexity often comes with computational challenges, akin to the trade-offs between importance sampling in particle filters and bootstrap filtering methods [60, 61]. To mitigate these issues, we have structured the method into several modular and easily testable steps. The two main approximations—using a neural network to learn time-shift distribution parameters and employing Laplace approximations—can both be replaced with exact computations (at the cost of efficiency) for validation. We recommend validating each component separately before applying the full framework. This includes ensuring that the neural network generalizes well beyond its training set, verifying the accuracy of time-shift distributions (e.g., by comparing against simulations from the stochastic model [17]), and using Laplace approximations only when justified (e.g., for normally distributed observation processes). Profile likelihoods can further confirm that the method accurately captures the shape of the exact likelihood, as demonstrated in Section 4 of S1 text..
@@
Another implementation consideration when applying to new models is that the calculation of time-shift distributions are model-specific. Our previous work [17] provides a package for computing the time-shift distributions given the branching process results that can be used for this. Additionally, the neural network must be redesigned and retrained for any new VL kinetics model as the underlying branching process structure is altered. While these challenges are not insurmountable, they are important factors when deciding whether to use this approach. However, our work here provides a general template for implementation making it substantially easier to extend the method to other models going forward.
Simulation-based validation of viral kinetics models has received relatively little attention in the literature, with most models being fitted directly to data and rarely tested on synthetic datasets. This is likely due to the challenge of generating biologically plausible simulations that capture realistic viral dynamics and experimental designs. In this work, we provide a preliminary validation study using multiple synthetic datasets generated from a single biologically plausible parameter set—drawn from literature estimates and consistent with our real dataset—and show that the true population parameters can be reasonably recovered. Validation and robustness of an inference method has many other dimensions, including model misspecification, varying sample sizes (both within and across individuals), population variability, and measurement error. Exploring inference performance across these factors is an important direction for future work, which would further assess method reliability and contribute to broader validation of viral kinetics models.
@@
A key advantage of our approach is its ability to model heterogeneity in viral trajectories, which is particularly relevant when initial infections involve a small number of cells. Beyond within-host modelling, incorporating stochasticity also has implications for between-host transmission studies [62, 63, 64, 65]. Small variations in VL trajectories can influence an individual’s infectiousness, affecting transmission dynamics in settings like household studies [39, 66, 67]. A more realistic representation of within-host variability could lead to improved estimates of key epidemiological parameters, which is an area of ongoing research.
@@
This paper introduces a methodological advancement for working with within-host viral dynamics (WHVDs). Rather than proposing a new model, we focus on developing computational tools that enable efficient Bayesian inference for existing stochastic VL kinetics models. Our interdisciplinary approach—combining applied mathematics, machine learning, and computational statistics—makes it feasible to fit these models to large datasets using biologically motivated, mechanistic, models. By leveraging modern computational techniques, we present a flexible framework that can accommodate the intricacies of viral dynamics and scale efficiently. A key enabler of this work is the use of the Julia programming language [42], which allows for seamless integration of model specification, automatic differentiation, and inference within a single environment. This capability makes it possible to conduct complex analyses on local hardware, reducing the reliance on dedicated high-performance computing clusters. Overall, the methods introduced here offer a powerful toolset for advancing the study of WHVDs and supporting future research into more realistic and scalable modelling frameworks.
@@
S1 text.
Supplementary text with additional methodological details. This text includes the derivation of the deterministic approximation and the key branching process results underlying the time-shift distributions. We provide a full derivation of the Laplace approximation and compare it with the exact calculation using profile likelihoods (S1 text. Fig. 1). Further details are given on the model and inference method, including the MCMC routine, as well as the construction of the neural network used for amortised optimisation.